OpenStudy (anonymous):

What is the equation of a line that passes through (4,7) and has a slope of -1/4 in point-slope form ?

OpenStudy (lucaz):

point slope form is: \[\large{ y-y_1=m(x-x_1) }\]

OpenStudy (anonymous):

it is y-7=-1/4(x-4)

OpenStudy (lucaz):

m is the slope, replace x1 by 4 and y1 by 7

OpenStudy (anonymous):

\[y=\frac{ x }{ 4 }+ 8\]

OpenStudy (anonymous):

(7)=(-1/4)*(4)+b

OpenStudy (anonymous):

no that is not point-slope form

OpenStudy (lucaz):

that is the y-intercept form \[\large{ y=mx+b}\]

OpenStudy (anonymous):

thats right @lucaz

OpenStudy (anonymous):

here are the choices \[y=\frac{ x }{ 4 }- \frac{ 7 }{ 4 } ..... y=- \frac{ x }{ 4}+\frac{ 7 }{ 4 }..... y=-\frac{ x }{ 4 }+8.....y=\frac{ x }{ 4 }+8\]

OpenStudy (anonymous):

@lucaz

OpenStudy (anonymous):

is the third one suppose to be x+8

OpenStudy (anonymous):

nope, that's how it was written bro

OpenStudy (anonymous):

what are your opinions guys ?

OpenStudy (lucaz):

doesnt make sense for me

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ 7}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -\cfrac{1}{4} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (anonymous):

my guess was the 3rd one lol

OpenStudy (anonymous):

I think its the third one but not certain. i got something else. sorry y=-1/4x+8

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ 7}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -\cfrac{1}{4} \\ \quad \\ y-{\color{blue}{ 7}}={\color{green}{ -\cfrac{1}{4}}}(x-{\color{red}{ 4}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \) distribute the -1/4, see what you get, and solve for

OpenStudy (lucaz):

geez, I inverted x and y.. from @jdoe0001 's post you see the answer is y = -x/4 + 8